# AI, Climate, and Regulation: From Data Centers to the AI Act

### v1

@@ 3.3 German 2023 Energy Efficiency Act
The Act broadens the scope of the reporting obligation to include even smaller data centers, upwards of 300 kW (Sec. 13). It also expands the duty to set up an energy management system to data centers and operators of ICT—i.e., customers of colocation data centers—of more than 50 kW (Sec. 12). Most importantly, it sets targets on energy efficiency and renewable energy use, requiring data centers to reach a PUE factor between 1.5 and 1.2 and an ERF of 10% to 20 % depending on their age (Sec. 11) and to run on 50 % renewable energy, increasing that factor to 100% by 1 Jan 2027 (Sec. 11). Lastly, it requires data center operators to inform their customers on an annual basis on the energy consumption directly attributable to them (Sec. 15).
While this implementation goes much further than the EED, it does not provide more detailed methodology. However, the German law may serve as a blueprint for further European regulation for efficiency and renewable energy targets. Especially the obligation to inform customers of their attributable energy consumption may prove an important prerequisite for AI developing companies to optimise energy consumption.
3.4 Bill in the U.S. Sentate for a 2024 AI Environmental Impacts Act
On 1 Feb 2024, the bill for an AI Environmental Impacts Act was introduced in the U.S. Senate by Senator Edward J. Markey (D-MA). [^12] US Senate, S.3732 - Artificial Intelligence Environmental Impacts Act of 2024. The bill was referred to the Committee on Commerce, Science and Transportation and there has not yet been a vote. [^13] Ibid. The future of the bill is therefore unclear.
Even if enacted, it would not contain any significant hard regulation. The bill orders the Environmental Protection Agency (EPA) to conduct a comprehensive study on AI’s environmental effects (Sec. 4) and requires the National Institute of Standards and Technology (NIST) to convene a consortium of stakeholders to identify the future measurements, methodologies, standards, and other appropriate needs to measure, report, and mitigate the environmental impacts of artificial intelligence (Sec. 5). It further requires NIST to develop a system for voluntary reporting of environmental impacts of artificial intelligence (Sec. 6). The EPA, NIST and the Department of Energy are to report to Congress on these three measures within four years (Sec. 7).
3.5 Interim Conclusion on Data Center Regulation
@@
One of the key strengths of the EU’s approach is the establishment of specific reporting obligations. Data center operators must collect and publicly report energy consumption, power utilization, water usage, waste heat utilization, and the use of renewable energy. These measures help create transparency and provide a foundation for future efficiency improvements. Additionally, the German implementation of the EED goes beyond mere reporting by setting specific targets for energy efficiency and renewable energy use in data centers, as well as requiring smaller data centers to comply. Germany’s approach might serve as a potential blueprint for broader EU regulation, particularly with its mandates to inform customers about their direct energy consumption, an essential factor for optimizing AI-related energy use.
@@
The situation in the U.S. contrasts with the EU approach. While a bill for an AI Environmental Impacts Act was introduced in the Senate, it remains unenacted and proposes only voluntary measures, such as studies and recommendations by federal agencies and institutions (EPA, NIST). The lack of binding rules for data centers in the U.S. means that AI’s environmental footprint could grow unchecked unless future legislation introduces more concrete regulatory measures.
4 AI Act: Gaps and Interpretation Challenges
@@
The AI Act applies further down in the value chain, targeting entities that develop and deploy AI systems. After introducing the key terminology used by the Act, we discuss the major climate-related obligations contained in it, with a particular focus on transparency, risk assessment and mitigation. As our analysis shows, the Act makes steps in the right direction, but fall short of a comprehensive regime for tackling climate-related risks of AI models.
4.1 Terminology of the AI Act: AI Models and Systems, Providers and Deployers
The AI Act only applies if an AI model or AI system is used by a specifically designated entity, such as a provider or deployer. The activity also needs to relate to the EU, typically either because the system is offered to persons in the EU or because its output is used in the EU (Art. 2(1) and Recital 22 AI Act).
@@ 4.1 Terminology of the AI Act: AI Models and Systems, Providers and Deployers
An AI model, though not explicitly defined in the AI Act, is understood to be a machine learning model that is trained on data and comprised of the parameters and the model architecture (Recital 104). AI models require the addition of further components, such as a user interface, to become AI systems (Recital 97). An AI system, in turn, is defined as a machine-based system ’designed to operate with varying levels of autonomy and that may exhibit adaptiveness after deployment, and that, for explicit or implicit objectives, infers, from the input it receives, how to generate outputs such as predictions, content, recommendations, or decisions that can influence physical or virtual environments’ (Art. 3(1)).
@@
Key actors in the AI value chain include providers and deployers. Providers develop or market AI models or systems (Art. 3(3)) while deployers use existing AI systems in a professional capacity (Art. 3(4)). Note that other operationally relevant actors, such as cloud service providers, data providers, data center operators, or other intermediaries are not directly subject to AI Act rules.
@@
The Act pays specific attention to general-purpose AI (GPAI) models, e.g., large language models (LLMs) like GPT-4, that are defined by the wide range of possible uses. Those with the most advanced capabilities, and therefore higher risk of negative effects, are labeled GPAI models with systemic risk (Art. 3(63)-(65)). An AI system that integrates a GPAI model is a general-purpose AI system (Art. 3(66)), e.g. ChatGPT. AI systems used in specific settings subject to product safety regulation, e.g., medical devices, and areas particularly sensitive for public safety or fundamental rights, e.g., law enforcement, migration, education, employment, credit scoring, health or life insurance, are called high-risk AI (HRAI) systems. These systems are subject to other rules, concerning training data, documentation, human in the loop and performance, inter alia (Art. 9-16). When GPAI systems are used in high-risk applications, rules may apply cumulatively (Recital 85, 97).
4.2 Provider or Deployer Status
@@
The AI Act contains energy- and climate-related transparency and risk management obligations, primarily for providers of AI systems or models. Under certain conditions, however, deployers turn into providers, which triggers much more onerous duties, also with respect to climate impacts. Hence, there is a palpable incentive for companies to avoid provider status.
@@
Concerning HRAI systems, pursuant to Art. 25(1), deployers can become providers if they market an existing HRAI system under their name or trademark, substantially modify an HRAI system or change the intended purpose of a non-high-risk AI system such that it falls under one of the high-risk sectors (e.g., employment; life- or health insurance; education). Specifically, if a deployer uses a GPAI system, such as ChatGPT, in a high-risk setting, for example by harnessing it for resume screening in hiring, they automatically assume the responsibilities of a provider of a HRAI system (Art. 25(1)(c)).
@@
An existing HRAI system can be altered. If the modification is substantial, the modifying entity becomes a new provider (Article 25(1)(b)). It is unclear, however, what precisely constitutes a "substantial modification" that would trigger provider status for the modifying entity. A formal interpretation suggests that any change to the model could be deemed substantial, based on Art. 3(23), which defines substantial modification as any unplanned change affecting compliance or intended purpose. On the other hand, a—more convincing—material interpretation would claim that provider status should only apply if the modification increases the model’s risk in a nonnegligible way. This aligns with the AI Act’s focus on risk mitigation, and also the language in Article 3(23) according to which any significant modification must affect compliance with the high-risk rules of the Act (or change the purpose).
@@
For GPAI models, Recitals 97, 109, and 111 suggest that any entity performing even minimal fine-tuning of a GPAI model is automatically classified as the provider of a new model, including all associated responsibilities [1]. For minor modifications this is disproportionate, even if Recital 109 states that obligations are limited to the fine-tuning itself [1]. A solution could be an analogy with Article 25(1)(b) so that provider status would only be triggered by a substantial modification made through fine-tuning, rather than any minor changes [1]. This is particularly important for SMEs using powerful models (GPAI models with systemic risk): if they become providers, they need to comply with the much more stringent obligations in Article 55 concerning evaluation and risk management, red teaming, incident reporting and cybersecurity.
4.3 Transparency
@@
The first climate-related obligation in the AI Act concerns transparency and reporting, but not with respect to data centers, but concerning providers of high-risk and GPAI models [1].
However, the rules come with significant ambiguities and loopholes. In the following, we detail the six most important ones.
@@
1) For high-risk AI systems, pursuant to Article 11(1) providers must document the computational resources utilized during the development, training, testing, and validation stages, as outlined in Annex IV(2) which does, however, not inlcude the energy consumption directly [1]. This omission hinders the capacity to evaluate and compare the environmental impact of such systems accurately. As a result, the environmental footprint must be estimated indirectly based on the recorded computational resources.
@@
2) The AI Act requires providers of general-purpose AI models to meet transparency obligations. Article 53(1)(a) mandates that providers maintain up-to-date technical documentation, including the details outlined in Annex XI. This annex requires the reporting of energy consumption, whether known or estimated, while estimates may be based on the computational resources used. However, this requirement exhibits a significant gap as it only covers the energy used during the model’s development phase, but leaves out the inference phase [1]. Recent research has shown that energy consumption during inference often exceeds that of the development phase very significantly over time [12, 21].
Although (pre-)training is the most noticeable driver of power consumption, the total energy usage by inferences becomes more significant in the long run when considering overall energy consumption. Eventually, the training of extensive LLMs will likely reach a point of saturation, though the specific timeline is uncertain and will depend on achieving adequate model performance or other factors like energy costs or data availability. Presently, convergence and subsequent training costs grow significantly with increasing model sizes. Nevertheless, once models are predominantly deployed and utilized in products, the energy consumption from inferences will rapidly surpass that of the initial model training phase. Although we are not aware of official numbers, this effect is likely to be already relevant for heavily utilized models such as those available by OpenAI’s ChatGPT.
In Luccioni et al. [12], the energy usage for inference in various large language models (Bloom) was assessed. They found that it takes between 200 and 500 million inferences, depending on the model’s size, for the total energy consumed by inferences to equate to the energy required for initial training. The necessity of incorporating inference energy consumption into mandatory reporting is underscored by Meta’s reported carbon emissions for various deployed AI models [21]. While (most of) these models are not LLMs, they demonstrate that the power consumption and associated carbon emmissions from inferences greatly exceeds that from training in production. There is a significant amount of ongoing research focused on optimizing the efficiency of training and inference for large-scale models. Implementing a reporting for inferences would create an incentive to balance not only performance and cost considerations but also explicitly focus on a reduced power consumption. Furthermore, obtaining information on the energy consumption of the cumulative inferences would additionally require the inclusion of model deployers and actually performed inference computations into the reporting obligations.
@@
To bridge this gap, we propose an alternative interpretation, based on the information provided to downstream actors (see also [1]). As a result of the AI Act, providers must supply downstream developers and relevant authorities with technical information on how to incorporate GPAI models into AI systems (Articles 53(1)(a) and (b), together with Annex XI and Annex XII). While energy consumption is not explicitly referenced in these provisions, they could reasonably be understood to include details about the computational resources required for inference. This would enable downstream providers to allocate resources more effectively and indirectly promote transparency around the energy consumption involved in each individual inference. However, to calculate the overall energy consumption for inferences, the total number of inferences would need to be available as well. While the typical energy costs of inferences should therefore be mandatorily disclosed by providers, deployers should be compelled to divulge information on the number of inferences, e.g., on an aggregate monthly level.
@@
3) Open-source (OS) GPAI models are largely excluded from transparency requirements unless they present a systemic risk, as outlined in Articles 2(12) and 53(2) [1]. The idea is that OS models, by definition, already disclose certain types of information. Hence, Recital 102 lists information on parameters, including weights, model architecture, and model usage as a prerequisite for systems to be considered OS. However, it does not mandate the disclosure of energy consumption. This results in an unwarrented lack of visibility concerning the environmental impact and energy usage of these models.
@@
4) Where the Act does mandate the disclosure of energy consumption, this information is restricted to authorities and is not accessible to downstream providers (unless the proposed interpretation from 2) is applied) or the general public, due to confidentiality clauses in Articles 21(3), 53(7), and 78(1) [1]. The limited availability of this data significantly reduces transparency and accountability, theryby weakening the potential for public oversight and market responses.
@@
5) The AI Act fails to address the greenhouse gas (GHG) emissions generated by AI applications, for instance in sectors like oil and gas exploration [7, 1]. For example, a recent investigation has revealed Microsoft’s aggressive pitch of its AI models to ExxonMobile to optimize fossil fuel exploration. [^14] Hao, Microsoft’s Hypocrisy on AI, The Atlantic, Sept. 13, 2024, https://www.theatlantic.com/technology/archive/2024/09/microsoft-ai-oil-contracts/679804/ Leaving such applications with significant climate effects out of scope creates a notable reporting gap.
@@
6) The Act also neglects to address the substantial water consumption, a key concern in data center operations [1]. While the Energy Efficiency Directive mandates reporting of water consumption for data centers, the AI Act requires no such reporting specifically for AI—as it does for energy use—nor does it cover operations outside the EU.
4.4 Environmental Risk Assessment and Mitigation

### v2

@@ 4.3. German 2023 Energy Efficiency Act
In Germany, the Energy Efficiency Act of 8 Nov 2023 implements the EED and establishes a national reporting scheme and additional requirements, including specific efficiency and renewable energy targets for data centers. The Act broadens the scope of the reporting obligation to include even smaller data centers, upwards of 300 kW (Sec. 13). It also expands the duty to set up an energy management system to data centers and operators of ICT—i.e., customers of colocation data centers—of more than 50 kW (Sec. 12). Most importantly, it sets targets on energy efficiency and renewable energy use, requiring data centers to reach a PUE factor between 1.5 and 1.2 and an ERF of 10% to 20 % depending on their age (Sec. 11), and to run on 50 % renewable energy, increasing that factor to 100% by 1 Jan 2027 (Sec. 11). Lastly, it requires data center operators to inform their customers on an annual basis on the energy consumption directly attributable to them (Sec. 15).
4.4. Regulation Outside the EU
Outside the EU, a number of countries is taking the lead by implementing concrete PUE goals, such as Singapore, Japan, China, and Australia, while the U.S. as one of the main contributors in the AI race shows a nuanced picture. However, it should be noted that some of these measures are limited to procurement policies or mere political commitments where it is unclear if there will be any penalty-enabled enforcement like in the German Energy Efficiency Act. A direct comparison of PUE target values should, therefore, be made with this caveat in mind.
Under the 2024 Green Data Center Roadmap, Singapore aims to reach a PUE of 1.3 or less within 10 years (Singapore Infocomm Media Development Authority, 2024). Japan, with its 2022 Energy Conservation Act, requires data center operators to take efficiency measures to reach a target PUE of 1.4 by 2030 (Business Network Editors, 2024). Meanwhile, in China, according to Li et al., policies issued from 2013 onward introduced efficiency targets, gradually decreasing the required PUE from 1.5 in 2013 to 1.3 in 2021 (Li et al., 2023a). However, the authors also admit that most recent data still shows a significant implementation gap with the actual PUE of most data centers ranging between 1.4 and 2.0. As part of Australia’s 2023 Net Zero in Government Operations Strategy, data center services procured by the government must demonstrate a five-star energy rating under the national NABERS scheme, including a PUE of 1.4 or lower. This requirement applies to procurement from members of the distinguished Data Centre Panel, and from 1 July 2025, it will extend to other companies and state-operated data centers (Department of Finance, Australian Government, 2023).
In the U.S., in the absence of federal efficiency requirements, the only direct regulation for data center efficiency is the California Green Building Action Plan 2015 implementing Executive Order B-18-12. Data centers that exceed a PUE of 1.5 are required to reduce their PUE by a minimum of 10 percent per year until they achieve a 1.5 or lower PUE. However, similar to the Australian action, this plan is only limited to state data centers (State of California, 2012; Department of General Services, 2014).
Another noteworthy strain of regulation is the climate reporting set forth in the SEC climate disclosure rules, and, on a state level, the California Climate Corporate Data Accountability Act, also known as SB 253. While the California Act goes further in including private companies and scope 3 indirect emissions, it is limited to companies with USD 1 B in annual revenue. For the SEC rules as well as SB 253 the focus is on the companies’ total emissions, therefore direct attribution to AI or computing energy consumption will be difficult.
Relating to AI more specifically, although not limited to data centers, is a bill for an AI Environmental Impacts Act that was introduced in the U.S. Senate by Senator Edward J. Markey (D-MA) on 1 Feb 2024 (U.S. Congress, 2024). The bill was referred to the Committee on Commerce, Science and Transportation, and has not yet been voted upon. Under the new administration, it is unlikely that the bill will pass Congress. However, even if enacted, it would not contain any significant hard regulation. The bill primarily mandates studies, stakeholder consultations, and voluntary reporting on AI’s environmental impacts without imposing significant regulatory obligations.
4.5. Discussion and Interim Conclusion on Data Center Regulation
@@
One of the key strengths of the EU’s approach is the establishment of specific reporting obligations. Data center operators must collect and publicly report energy consumption, power utilization, water usage, waste heat utilization, and the use of renewable energy. These measures help create transparency and provide a foundation for future efficiency improvements. Additionally, the German implementation of the EED goes beyond mere reporting by setting specific targets for energy efficiency and renewable energy use in data centers, as well as requiring smaller data centers to comply. Germany’s approach might serve as a potential blueprint for broader EU regulation, particularly mandating data center operators to inform customers about their direct energy consumption, an essential factor for optimizing AI-related energy use.
@@
In contrast, the situation outside the EU shows a very diverse landscape, ranging from government procurement to industry wide political commitments and concrete legal requirements with varying implementation success.
5. AI Act: Gaps and Interpretation Challenges
@@
The AI Act applies further down in the value chain, targeting entities that develop and deploy AI systems. After introducing the key terminology used by the Act, we discuss the major climate-related obligations contained in it, with a particular focus on transparency, risk assessment and mitigation. As our analysis shows, the Act makes steps in the right direction, but falls short of a comprehensive regime for tackling climate-related risks of AI models.
5.1. Terminology of the AI Act: AI, Providers, and Deployers
@@
The AI Act only applies if an AI model or AI system is used by a specifically designated entity, such as a provider or deployer. The activity also needs to relate to the EU, typically either because the system is offered to persons in the EU or because its output is used in the EU (Art. 2(1) and Recital 22 AI Act). An AI system is defined as a machine-based system ’designed to operate with varying levels of autonomy and that may exhibit adaptiveness after deployment, and that, for explicit or implicit objectives, infers, from the input it receives, how to generate outputs such as predictions, content, recommendations, or decisions that can influence physical or virtual environments’ (Art. 3(1)).
@@
Key actors in the AI value chain include providers and deployers. Providers develop or market AI models or systems (Art. 3(3)) while deployers use existing AI systems in a professional capacity (Art. 3(4)). Note that other operationally relevant actors, such as cloud service providers, data providers, data center operators, or other intermediaries, are not directly subject to AI Act rules.
@@
The Act pays specific attention to general-purpose AI (GPAI) models, e.g., large language models (LLMs) like GPT-4, that are defined by the wide range of possible uses. Those with the most advanced capabilities, and therefore higher risk of negative effects, are labeled GPAI models with systemic risk (Art. 3(63)-(65)). An AI system that integrates a GPAI model is a general-purpose AI system (Art. 3(66)), e.g., ChatGPT. AI systems used in specific settings subject to product safety regulation, e.g., medical devices, and areas particularly sensitive for public safety or fundamental rights, e.g., law enforcement, education, employment, or credit scoring, are called high-risk AI (HRAI) systems. These systems are subject to other rules, concerning training data, documentation, human in the loop, and performance, inter alia (Art. 9-16). When GPAI systems are used in high-risk applications, rules may apply cumulatively (Recital 85, 97).
@@
The AI Act contains energy- and climate-related transparency and risk management obligations, primarily for providers of AI systems or models. Under certain conditions (e.g., fine-tuning), however, deployers become providers, which triggers much more onerous duties, also with respect to climate impacts (Art. 25(1), Recital 109). Hence, there is a palpable incentive for companies to avoid provider status.
@@
Concerning HRAI systems, pursuant to Art. 25(1), deployers can become providers if they market an existing HRAI system under their name or trademark, substantially modify an HRAI system, or change the intended purpose of a non-high-risk AI system such that it falls under one of the high-risk sectors (e.g., employment; life- or health insurance; education). Specifically, if a deployer uses a GPAI system, such as ChatGPT, in a high-risk setting, for example, by harnessing it for resume screening in hiring, they automatically assume the responsibilities of a provider of an HRAI system (Art. 25(1)(c)).
@@
An existing HRAI system can be altered. If the modification is substantial, the modifying entity becomes a new provider (Article 25(1)(b)). It is unclear, however, what precisely constitutes a ”substantial modification” that would trigger provider status for the modifying entity. A formal interpretation suggests that any change to the model could be deemed substantial, based on Art. 3(23), which defines substantial modification as any unplanned change affecting compliance or intended purpose. On the other hand, a—more convincing—material interpretation would claim that provider status should only apply if the modification increases the model’s risk in a nonnegligible way. This aligns with the AI Act’s focus on risk mitigation and also the language in Article 3(23) according to which any significant modification must affect compliance with the high-risk rules of the Act (or change the purpose).
@@
For GPAI models, Recitals 97, 109, and 111 suggest that any entity performing even minimal fine-tuning of a GPAI model is automatically classified as the provider of a new model, including all associated responsibilities (Anonymous, 2024a). For minor modifications this is disproportionate, even if Recital 109 states that obligations are limited to the fine-tuning itself (Anonymous, 2024a). A solution could be an analogy with Article 25(1)(b) so that provider status would only be triggered by a substantial modification made through fine-tuning, rather than any minor changes (Anonymous, 2024a).
This is particularly important for SMEs using powerful models (GPAI models with systemic risk): if they become providers, they need to comply with the much more stringent obligations in Article 55 concerning evaluation and risk management, including environmental risk.
5.2. Legislative History and Intent
In the original Commission proposal of April 2021, the AI Act only included minimal references to environmental effects, mostly as a potential benefit to climate change mitigation (see above, 2.), and in the section on voluntary commitments (Pagallo et al., 2022). The General Approach adopted by the Council in December 2022 continued this narrow framing. However, the momentum changed, particularly with the position of the European Parliament (EP), published in June 2023. The suggested EP sustainability amendments to the AI Act introduce environmental principles for AI development, preferential funding for eco-friendly AI, mandatory energy and resource consumption reporting, sustainability risk assessments (SIAs) for foundation models, and Commission-led guidance on environmental impact measurement and review (Hacker, 2024b). During the final trilogue negotiations, several of these amendments were dropped (e.g., the principle; explicit SIAs). However, simultaneously, the impact of fundamental rights was strengthened, and environmental protection listed as one such right that must be considered in a range of provisions (Hacker, 2024b; Pagallo et al., 2022). This is supported by the mention of environmental protection as a key policy goal of the AI Act in Article 1 and Recitals 1, 2 and 8.
5.3. Transparency
@@
The first climate-related obligation in the AI Act concerns transparency and reporting, however not with respect to data centers, but concerning providers of high-risk and GPAI models (Anonymous, 2024a). Yet, the rules come with significant ambiguities and loopholes. In the following, we detail the six most important ones.
@@
1) For high-risk AI systems, pursuant to Article 11(1) providers must document the computational resources utilized during the development, training, testing, and validation stages, as outlined in Annex IV(2) which does, however, not include the energy consumption directly (Anonymous, 2024a). This omission hinders the capacity to evaluate and compare the environmental impact of such systems accurately. As a result, the environmental footprint must be estimated indirectly based on the recorded computational resources.
@@
2) The AI Act requires providers of general-purpose AI models to meet transparency obligations. Article 53(1)(a) mandates that providers maintain up-to-date technical documentation, including the details outlined in Annex XI. This annex requires the reporting of energy consumption, whether known or estimated, while estimates may be based on the computational resources used. However, this requirement exhibits a significant gap as it only covers the energy used during the model’s development phase, but leaves out the inference phase (Anonymous, 2024a). Recent research has shown that energy consumption during inference often exceeds that of the development phase very significantly over time (Luccioni et al., 2024b; Wu et al., 2022).
@@
To bridge this gap, we propose an alternative interpretation, based on the information provided to downstream actors (see also (Anonymous, 2024a)). As a result of the AI Act, providers must supply downstream developers and relevant authorities with technical information on how to incorporate GPAI models into AI systems (Articles 53(1)(a) and (b), together with Annex XI and Annex XII). While energy consumption is not explicitly referenced in these provisions, they could reasonably be understood to include details about the computational resources required for inference because downstream providers need this information to properly allocate resources, e.g., purchasing computing power for operating the model. As a side-effect, this information could be used to calculate energy consumption for individual inferences. However, to calculate the overall energy consumption for inferences, the total number of inferences would need to be available as well. While the typical energy costs of inferences should therefore be mandatorily disclosed by providers, deployers should be compelled to divulge information on the number of inferences, as well, e.g., on an aggregate monthly level.
@@
3) Open-source (OS) GPAI models are largely excluded from transparency requirements unless they present a systemic risk, as outlined in Articles 2(12) and 53(2) (Anonymous, 2024a). The idea is that OS models, by definition, already disclose certain types of information. Hence, Recital 102 lists information on parameters, including weights, model architecture, and model usage as a prerequisite for systems to be considered OS. However, it does not mandate the disclosure of energy consumption. This results in an unwarrented lack of visibility concerning the environmental impact and energy usage of these models.
@@
4) Where the Act does mandate the disclosure of energy consumption, this information is restricted to authorities and is not accessible to downstream providers (unless the proposed interpretation from 2) is applied) or the general public, due to confidentiality clauses in Articles 21(3), 53(7), and 78(1) (Anonymous, 2024a). The limited availability of this data significantly reduces transparency and accountability, theryby weakening the potential for public oversight and market responses.
@@
5) The AI Act fails to address the greenhouse gas (GHG) emissions generated by AI applications, for instance in sectors like oil and gas exploration (Kaack et al., 2022; Anonymous, 2024a). For example, a recent investigation has revealed Microsoft’s aggressive pitch of its AI models to ExxonMobile to optimize fossil fuel exploration (Hao, 2024). Leaving such applications with significant climate effects out of scope creates a notable reporting gap.
@@
6) The Act also neglects to address the substantial water consumption, a key concern in data center operations (Anonymous, 2024a). While the Energy Efficiency Directive mandates reporting of water consumption for data centers, the AI Act requires no such reporting specifically for AI—as it does for energy use—nor does it cover operations outside the EU.
5.4. Environmental Risk Assessment and Mitigation